# FRASA: An End-to-End Reinforcement Learning Agent for Fall Recovery and Stand Up of Humanoid Robots

### v1

@@ I Introduction
This paper does not focus on the problem of push recovery but rather on fall recovery, which involves creating new points of contact with the ground as a strategy to regain stability in minimal time. The approach usually used in robotics competitions [2] [3] [4] is to wait for the fall to end before starting a stand up routine, generally based on key frame animations [17]. This method is simple to implement but lacks versatility, requires extensive fine-tuning, and must be frequently adjusted due to motor wear and loss of precision.
@@
Tests are conducted on Sigmaban humanoid robots [21] to compare FRASA with a Key Frame Based (KFB) approach. The KFB method used for comparison was deployed in the RoboCup 2023 by the Rhoban Team, world champion of the KidSize League. Real robots experiments are presented in the following video [^1] https://youtu.be/NL65XW0O0mk.
@@ IV Method
In this paper, we present FRASA: an end-to-end DRL approach for humanoid robot autonomous stand up and fall recovery from external disturbances. The agent is trained in a full-body physics simulator [26], using a robot’s detailed collision model. In the simulation, the robot’s degrees of freedom (DoFs) are position controlled to reflect the servomotors behavior.
@@ IV-A Environment initialization
Figure 2: Left: target posture for the recovery movement. Right: state vector components, including the trunk pitch $\theta$ and the 5 explicited DoFs.
@@ IV-D Reward
Desired state proximity reward $R_{\text{state}}=\exp(-w_{1}\times\|\psi_{t}-\psi_{\text{desired}}\|^{2})$
@@
The weights $0\leq w_{2},w_{3}\ll 1$ are small penalizations intended to smooth the learning process and prevent the robot from adopting mechanically feasible but undesirable strategies.
When the robot reaches the desired position, the reward is approximately 1, providing a clear unit of measure for the agent’s performance in terms of maintaining the upright posture over time.
@@ V-B Physical Simulation and Modeling
The physics engine used to simulate the environment during training is MuJoCo[26]. A precise CAD model of the robot is exported to MuJoCo[30], ensuring accurate geometry, mass distribution, and joint limits, closely matching the real robot.
@@ V-E Stand Up Experiment
A first experiment is conducted to compare the ability to stand up from prone and supine positions using both a KFB recovery process and FRASA. A KFB stand up method defines arbitrary key positions determined thanks to expert knowledge for the robot to pass through, interpolating motor positions between them to create a stand up trajectory. The KFB method used in the experiments was developed for RoboCup 2023 by the Rhoban Team and demonstrated both its effectiveness and efficiency [28].
The robot is placed on its back and face down, aiming to stand and begin walking. The metric studied is the time from movement initiation to the start of walking. To initiate walking, the error in the state vector components—comprising the 5 controlled joints and the trunk pitch—must remain below 5° for 0.5 seconds.
@@ V-F Fall Recovery Experiment
In this experiment, the robot is placed under the pendulum and subjected to repeated disturbances with varying impact intensities. After a disturbance, the robot waits to return to a stable position close to the target posture (cf. Fig. 2) before beginning to walk in place.
The metric studied is the duration of disturbance rejection. Time measurement begins when any value in the state vector deviates by more than 5° from the neutral walking posture values. Once in this unstable state, all values in the state vector must have an error of less than 5° relative to the target posture for $0.5$ s for the robot to be considered stable again and able to initiate walking. The time measurement ends at the beginning of the walking phase.
@@ VII Conclusion
The end-to-end nature of FRASA allows it to adapt to various disturbance intensities without requiring expert tuning. However, while leveraging the humanoid robot’s symmetry accelerates training, it also limits the versatility of the emerging behaviors, particularly in handling lateral disturbances.

### v2

@@ I Introduction
This paper does not focus on the problem of push recovery but rather on fall recovery, which involves creating new points of contact with the ground as a strategy to regain stability in minimal time. This approach is crucial when external forces are too strong on the robot to prevent a fall. In robotics competitions [2][3][4], fall recovery is usually handled by allowing the fall to complete before initiating a stand-up routine, which is often based on key frame animations [17]. This method is simple to implement but lacks versatility, requires extensive fine-tuning, and must be frequently adjusted due to motor wear and loss of precision.
@@
Tests are conducted on Sigmaban humanoid robots [21] to compare FRASA with a Key Frame Based (KFB) approach [17]. The KFB method used for comparison was deployed in the RoboCup 2023 by the Rhoban Team, world champion of the KidSize League. Real robots experiments are presented in the following video [^1] https://youtu.be/NL65XW0O0mk and the open-source code for this project is available on our GitHub repository [^2] https://github.com/Rhoban/frasa for further use and contributions.
@@ IV Method
In this paper, we present FRASA: an end-to-end DRL approach for humanoid robot autonomous stand up and fall recovery from external disturbances. The agent is trained in a full-body physics simulator [26], incorporating a robot’s detailed collision model. In simulation, the robot’s degrees of freedom (DoFs) are position controlled to reflect servomotors behavior. The core contribution of FRASA lies in formulating an environment that enables efficient learning with actor-critic algorithms, such as SAC and, more recently, CrossQ.
@@ IV-A Environment initialization
Figure 2: Left: target posture for the recovery movement. Right: pose vector $\psi_{target}$ components, including the trunk pitch $\theta$ and the 5 explicited DoFs.
@@ IV-D Reward
Desired state proximity reward $R_{\text{state}}=\exp(-w_{1}\times\|\psi_{t}-\psi_{\text{target}}\|^{2})$
@@
The weights $0\leq w_{2},w_{3}\ll 1$ are small penalizations intended to smooth learning and prevent the robot from adopting mechanically feasible but undesirable strategies.
When the robot reaches the desired position, the reward is approximately 1 thanks to $R_{\text{state}}$, providing a clear unit of measure for the agent’s performance in terms of maintaining upright posture over time.
@@ V-B Physical Simulation and Modeling
The physics engine used for training is MuJoCo[26]. A precise CAD model of the robot is exported to MuJoCo[30], ensuring accurate geometry, mass distribution, and joint limits, closely matching the real robot.
@@ V-E Stand Up Experiment
A first experiment is conducted to compare the ability to stand up from prone and supine positions using both a KFB recovery process and FRASA. A KFB stand up method defines arbitrary key positions determined thanks to expert knowledge for the robot to pass through, interpolating motor positions between them to create a stand up trajectory [17]. The KFB method used in the experiments was developed for RoboCup 2023 by the Rhoban Team and demonstrated both its effectiveness and efficiency [28].
The robot is placed on its back, face down, aiming to stand and begin walking. The metric studied is the time from movement initiation to the start of walking. To initiate walking, the error in the state vector components—comprising the 5 controlled joints and the trunk pitch—must remain below 5° for 0.5 seconds.
@@ V-F Fall Recovery Experiment
In this experiment, the robot is placed under the pendulum and subjected to repeated disturbances with varying impact intensities. After a disturbance, the robot waits to return to a stable position close to the target posture (cf. Fig. 2) before starting to walk in place.
The metric studied is the duration of disturbance rejection. Time measurement begins when any value in the state vector deviates by more than 5° from the neutral walking posture values. Once in this unstable state, all values in the state vector must have an error of less than 5° relative to the target posture for 0.5 seconds for the robot to be considered stable again and able to initiate walking. The time measurement ends at the beginning of the walking phase.
@@ VII Conclusion
The end-to-end nature of FRASA allows it to adapt to various disturbance intensities without requiring expert tuning. However, while leveraging the humanoid robot’s symmetry accelerates training, it also limits the versatility of emerging behaviors, particularly in handling lateral disturbances.